### v1

# MATHGLANCE: Multimodal Large Language Models Do Not Know Where to Look in Mathematical Diagrams

Abstract
Diagrams serve as a fundamental form of visual language, representing complex concepts and their interrelationships through structured symbols, shapes, and spatial arrangements. Unlike natural images, their inherently symbolic and abstract nature poses significant challenges for Multimodal Large Language Models (MLLMs). However, current benchmarks conflate perceptual and reasoning tasks, making it difficult to assess whether MLLMs genuinely understand mathematical diagrams beyond superficial pattern recognition. To address this gap, we introduce MATHGLANCE, a benchmark specifically designed to isolate and evaluate mathematical perception in MLLMs. MATHGLANCE comprises 1.2K images and 1.6K carefully curated questions spanning four perception tasks: shape classification, object counting, relationship identification, and object grounding, covering diverse domains including plane geometry, solid geometry, and graphical representations. Our evaluation of MLLMs reveals that their ability to understand diagrams is notably limited, particularly in fine-grained grounding tasks. In response, we construct GeoPeP, a perception-oriented dataset of 200K structured geometry image-text pairs explicitly annotated with geometric primitives and precise spatial relationships. Training MLLM on GeoPeP leads to significant gains in perceptual accuracy, which in turn substantially improves mathematical reasoning. Our benchmark and dataset establish critical standards for evaluating and advancing multimodal mathematical understanding, providing valuable resources and insights to foster future MLLM research.
Figure 1: Performance comparison of current Multimodal Large Language Models (MLLMs) on MATHGLANCE reveals a stark contrast: while humans can solve fundamental perception tasks “at a glance”, these tasks remain highly challenging for MLLMs—particularly in fine-grained grounding. SVE-Math-DeepSeek $\!\!{}^{+}$-7B, trained with the proposed perception-oriented GeoPeP, shows significantly enhanced perceptual capabilities.
1 Introduction
Computer vision research has traditionally focused on developing models that can perform fundamental perception tasks [17, 30] such as object detection [36, 56], segmentation [28, 37], and spatial reasoning [12]. These capabilities form the foundation for higher-level visual reasoning and decision-making. However, recent advances in Multimodal Large Language Models (MLLMs) [1, 32, 42] have shifted the paradigm toward vision-language integration, where visual information is processed through natural language interfaces rather than specialized perception modules.
Figure 2: Illustration of diagram-caption-based alignment training datasets, w.r.t., AutoGeo, MAVIS, and our proposed GeoPeP (Fig. 2(a)). Fig. 2(b) demonstrates a positive correlation between low-level perception and high-level reasoning tasks, evaluated on MATHGLANCE and MathVista. Clear diagram perception leads to substantial improvements in mathematical reasoning performance.
Despite their remarkable successes in general vision tasks, current MLLMs face considerable challenges in accurately interpreting mathematical diagrams—abstract visual representations characterized by precise geometric structures and symbolic notations [34, 53]. These diagrams provide a unique perceptual challenge: the explicit structural constraints require genuine visual interpretation rather than pattern recognition or textual heuristics, positioning them as an essential benchmark for evaluating the true visual perception capabilities of multimodal models. Such diagrams are pervasive across educational contexts, scientific communication, and problem-solving in STEM disciplines. Developing models capable of understanding symbolic information in diagrams, is a critical milestone in advancing machine intelligence  [14, 13].
Recent benchmarks like MathVista [34] and MathVerse [53] have attempted to evaluate mathematical visual reasoning in MLLMs. However, these works conflate perception with reasoning by assessment on complex reasoning tasks that combine diagram interpretation with numerical computation and proof generation. Thus, it remains unclear whether the performance truly reflects the models’ ability to comprehensively understand the symbolic information in diagrams. Misinterpretations at the perception level propagate errors downstream, resulting in incorrect reasoning outcomes and frequent model hallucinations [7, 44, 23].
To address this critical gap, we design a benchmark to isolate and rigorously evaluate mathematical perception in MLLMs. Our MATHGLANCE features problems that humans can solve “at a glance” without extensive reasoning. The benchmark contains 1,198 images and 1,609 carefully designed perception-oriented questions across four distinct task categories: shape classification, object counting, relationship identification, and object grounding. These tasks span three core mathematical domains—plane geometry, solid geometry, and graphical representations (line, bar, and pie graphs)—and question formats (multiple-choice, true/false, and free-form).
Figure 3: Sampled MATHGLANCE  examples from plane geometry w.r.t. each question-answer (Q&A) type, covering shape classification, object counting, relationship identification, and object grounding (from left to right). The green dotted bounding boxe is shown for illustration purposes only and are not provided as input to the models.
Through comprehensive evaluation of current MLLMs on MATHGLANCE, we explore the following questions:
  1. 1.
    Do current MLLMs genuinely perceive mathematical diagrams? Fig. 1 presents the performance of eight current MLLMs—comprising six generic models (both open-source and proprietary) and two math-specific model—across four perception-focused tasks in MATHGLANCE (see Tab. 2 for more results). Among the generic MLLMs, Qwen2.5-VL-7B [6] achieves the highest average accuracy, followed by GPT-4o. But, those models exhibit severe hallucination issues, often responding to simple questions with unnecessarily long chain-of-thought (CoT) reasoning and generating irrelevant image content (see Fig.6). The math-specific SVE-Math-DeepSeek [55], which incorporates a domain-specific visual encoder trained on multi-task geometric primitive detection, achieves comparable performance to Qwen2.5-VL-7B on shape classification and relationship identification tasks. Nevertheless, all evaluated models struggle significantly with fine-grained bounding box grounding. Even those trained on large-scale grounding annotations with nearly 1T natural images—such as Qwen2.5-VL-7B and DeepSeek-VL2-Small—achieve less than 20% accuracy on this task.
    To alleviate this issue, we propose GeoPeP, a Geomatic PercePtion-oriented dataset. Unlike existing image-caption-based visual training datasets (e.g., MAVIS [54], AutoGeo [21]), GeoPeP provides explicit structured diagram information, including shapes, attributes, locations, and relationships (see Fig.2(a)). We first train only the vision-language projector of SVE-Math-DeepSeek using this alignment dataset. Then, we construct instruction conversational data for a second self-fine-tuning (SFT) stage, jointly updating both the projector and the large language model. The resulting model, SVE-Math-DeepSeek $\!{}^{+}$, achieves the highest performance among both generic and math-specific models, particularly excelling in grounding tasks with a +79% improvement.
  2. 2.
    Does stronger perceptual ability lead to better reasoning performance? We hypothesize that the more a model can “see”, the more it can “reason” —that is, clear diagram perception would support stronger reasoning ability. As shown in Fig.2(b), training with our proposed alignment dataset significantly improves perception performance compared to AutoGeo and MAVIS (by 23.0% and 35.6%, respectively). In contrast, the other two variants perform worse than the baseline (SVE-Math-DeepSeek-7B), particularly with MAVIS diagram-caption pairs, due to high uncertainty and out-of-distribution nature of real-world geometric diagrams.
    For a fair comparison, we use the same reasoning training dataset (MathV360K [41]) and apply our instruction-style perception sampels to further fine-tune each model. Our model achieves 28.1% accuracy on the MathVista reasoning benchmark—an improvement of $\sim$4over others. This gain is non-trivial, considering that Multi-Math-7B [38], trained on 600K samples (300K image captions+300K CoT rationales) with additional reinforcement learning, achieves only 26.9% accuracy. Our model accurately identifies relevant visual elements, enabling it to generate more valid and faithful reasoning steps (see model responses in Figs. 7/15-16).
In summary, our contributions are as follows:
  • We introduce MATHGLANCE, a comprehensive benchmark for evaluating mathematical perception across geometry and graphical domains. MATHGLANCE isolates perception capabilities through carefully designed tasks and the quality of the data is ensured via rigorous human verification.
  • We provide the systematic investigation of how perception capabilities influence mathematical reasoning, quantitatively analyzing MLLMs’ perceptual sensitivity to various factors and establishing the correlation between perception accuracy and reasoning performance (see §3.3).
  • We construct GeoPeP, a Geomatic PercePtion-oriented dataset comprising 200K high-quality geometry image-text pairs (100K diagram-caption + 100K conversation) that explicitly struct diagrams into shapes, attributes, locations and relationships. GeoPeP provides precise visual grounding cues and minimizes ambiguities common in existing caption-based datasets, which can significantly improve MLLMs’ visual perception accuracy and downstream reasoning performance.
We envision MATHGLANCE as a valuable testbed to bridge the gap between low-level visual perception and high-level mathematical reasoning, ultimately facilitating the development of more robust and reliable multimodal models for mathematical understanding.
2 MATHGLANCE
This section first introduces MATHGLANCE (§2.1) with a data distribution analysis. We then describe the synthetic construction process for plane geometry, as well as the reformatted CLEVR [24] and FigureQA [25] datasets for solid geometry and graphs in §2.2. Finally, we introduce the perception-oriented GeoPeP, and highlight its unique characteristics in comparison to existing mathematical visual alignment training datasets (§2.3).
Table 1: Key statistics of MATHGLANCE are summarized in Tab. 3(a), and the subject-task distribution is illustrated in Fig. 3(b).
        Statistic         Number
        Total questions         1,609
         - Multiple-choice questions         893 (55.5%)
         - Free-form questions         406 (25.2%)
         - True-false questions         310 (19.3%)
        Unique number of images         1,198
        Unique number of questions         1,514
        Unique number of answers         380
        Total questions         1609
         - Classification questions         489 (30.4%)
         - Counting questions         401 (24.9%)
         - Relation questions         313 (19.5%)
         - Grounding questions         406 (25.2%)
        Maximum question length         200
        Maximum answer length         4
        Maximum choice number         4
        Average question length         118.0
        Average answer length         1.8
        Average choice number         3.5
        Average question length         16.09
        Average answer length         1.21
        Average choice number         3.40
2.1 Overview
Existing mathematical visual reasoning benchmarks often conflate perception with higher-level reasoning tasks, such as numeric calculations and proof generation. We introduce MATHGLANCE, a novel benchmark designed to evaluate MLLMs’ mathematical perception-demanding abilities through both quantitative and qualitative analysis across coarse-to-fine granularity levels.
Data Composition. Detailed statistics for data composition are presented in Tab. 3(b). To comprehensively assess the perceptual abilities of MLLMs in mathematical contexts, our evaluation images cover plane geometry (66%), solid geometry (20%), and graphs (14%), including lines, bars, and pie charts. To facilitate MLLM evaluation, we formulate all tasks—except for bounding box grounding (free-form)—as multiple-choice or true/false question-answering problems. In total, we contribute 1,609 questions and 1,198 unique images, ensuring an even distribution across the different perception tasks.
Categorization. The benchmark encompasses shape classification, object counting, relationship identification, and object grounding. Fig. 3 shows example illustrations of plane geometry (see §B for additional examples). Key features are as follows:
  • Shape classification is a classic vision task where the model identifies an object’s class based on its attributes, i.e., vertices, material, color, and size. Our dataset includes a diverse set of geometric categories, comprising 16 basic shapes for plane geometry, 3 CLEVR-defined [24] objects for solid geometry, and 5 graphical elements as defined in FigureQA [25].
  • Object counting requires models to determine either the total number of objects in an image or a specific shape count, i.e., the number of circles or triangles present.
  • Relationship identification evaluates models’ understanding of 4 spatial and over 10 mathematical relationships between pairs of geometric primitives.
  • Object grounding evaluates fine-grained localization by requiring MLLMs to accurately predict the top-left and bottom-right coordinates in the format (x1, y1, x2, y2) for an object within the image. This ensures that models can precisely identify and localize geometric structures based on textual descriptions.
These four tasks are designed to isolate and assess fundamental perception capabilities, distinguishing low-level visual recognition from high-level reasoning.
Figure 4: The synthetic construction process for plane geometry. We synthesize geometric figures by randomly sampling elements from the geometric shape pool and relationship pool, ensuring consistency through a verifier that enforces logical constraints based on manually designed rules, fundamental mathematical principles, and prerequisite points. All visual elements are structured and saved in JSON format. Images are rendered using the Matplotlib package, and corresponding Q&A pairs are generated using a template-based pipeline.
2.2 Dataset Construction
2.2.1 Synthetic Data Engine for Plane Geometry
We design a data generation engine to synthesize real-world geometric images along with structured JSON annotations. This pipeline enables controlled shape generation, relationship modeling, and visual attribute assignment, ensuring diverse and well-balanced datasets across perception tasks.
Structured Annotations for Geometric Primitives. Fig. 4 describes the entire generation process. Inspired by AlphaGeometry [43], we use geometric clauses as fundamental units to construct complex plan geometric figures. A geometric clause is a formalized description of basic geometric objects and mathematical relationships, along with their properties or attibutes, a.k.a. prerequisite points. We first construct two geometry substrate pools, one containing 16 different geometric shapes (i.e., scalene triangle, isosceles triangle, square, rectangle, parallelogram, isosceles trapezoid, right trapezoid, pentagon, circle $\dots$ ellipse and segment), and the other defining 10 mathematical relationships (i.e., on, intersection, parallel, perpendicular, tangent, incircle, $\dots$ reflection and angle bisector). We then randomly sample one or more substrates from these pools and pass them through a verifier, which ensures that logically paired shapes and relationships are preserved in the construction of valid geometric images. The verifier makes decisions based on either manually designed rules, fundamental mathematical knowledge, or prerequisite points. For example, parallel lines cannot intersect, making an intersection clause invalid. Similarly, angle trisection requires a predefined angle, which cannot be constructed without enough prerequisite points for angle consturction. In general, the chosen relationships are enforced by introducing additional shapes into the figure, ensuring that the specified relationships are accurately maintained and geometrically consistent throughout the construction process. We finally save the outputs as structured JSON annotations for image rendering using the Matplotlib package [22] and generating question-answer pairs via template-based pipelines.
Image Generation. The JSON annotations define foreground and background styles, i.e., colors sampled from a monochromatic palette, line width, and font size, as well as object shape information (class names), attributes (vertices labeled with random letters), bounding box locations, and mathematical relationships. Notably, spatial relationships are generated based on the bounding box locations of two objects (e.g., top-left, top-right, bottom-left, bottom-right). Similar to the image rendering process in  [43], we translate geometric clauses into visual representations using Python code. This process first determines the coordinates of each point defined in the clause for basic shape visualization, and then new primitives are generated based on the specified relationships. Additionally, we apply image augmentations to increase task difficulty, e.g., adding Gaussian noise, drawing irregular scribbles, and incorporating wedge-shaped symbols for congruent angles or auxiliary lines (see Figs. 9 and 11 in the Appendix).
Question-Answer Generation. Based on the structured annotations, we employ a template-based pipeline to generate multiple-choice and true-false questions across four perception tasks. For example, in constructing Q&A pairs for shape classification, we first randomly select an object and its associated attributes from the given figure. We then formulate a question by filling placeholders in our carefully designed templates. For true-false questions, the ground truth answer is directly derived from the JSON annotations. For multiple-choice questions, we generate plausible distractors to challenge the model’s geometric perception, combining them with the correct answer into a set of four choices. Distractors are selected from geometric candidate pools, ranging from visually similar (e.g., equilateral vs. isosceles triangles) to visually dissimilar pairs (e.g., equilateral triangle vs. circle). Further details on the Q&A generation templates are provided in Tab. 7 of the Appendix.
2.2.2 Reformat Dataset for Solid Geometry & Graphs
We leverage the public datasets CLEVR [24] and FigureQA [25] to construct benchmarks for solid geometry and graphs, respectively. Since their original annotation formats are incompatible with our template-based Q&A generation pipeline, we reformat them into structured JSON, consistent with the format used in our synthetic data process. Once reformatted, we apply the same Q&A generation strategy as used for plane geometry. See §D for details.
After generating our synthetic data and collecting public datasets, we conduct a comprehensive review to verify answer accuracy, ensure consistency between questions and diagrams, and confirm relevance to the four perception tasks, ensuring high-quality and precise dataset annotations.
Table 2: Performance comparison of different MLLMs on MATHGLANCE across plane geometry, solid geometry, and graphs. cls, cnt, grd, and rlat represent different question categories: shape classification, object counting, object grounding, and relationship identification, respectively. all indicates the overall accuracy, calculated as the ratio of correctly answered questions to the total number of questions in the benchmark, while Avg. denotes the average all score across all subjects.
Model Size Avg. Plane Geometry Soild Geometry Graphs
all cls cnt grd rlat all cls cnt grd rlat all cls cnt grd rlat
Open-Source Generic MLLMs
LLaVA-v1.5 [32] 7B 33.3 29.2 29.0 39.6 14.2 37.5 31.6 43.0 42.3 0.0 31.3 39.0 76.8 35.2 0.0 39.4
LLaVA-v1.5 [32] 13B 35.4 32.8 29.3 40.4 23.5 42.0 35.9 60.5 38.1 0.0 35.0 37.6 63.8 42.6 0.0 45.5
mPLUG-owl3 [48] 7B 50.0 36.4 46.7 41.6 3.9 58.5 65.3 95.4 83.5 0.0 62.5 48.2 59.4 77.8 0.0 66.7
InternLM-XComposer2 [16] 7B 55.6 35.8 49.4 48.8 0.0 47.0 62.9 90.7 86.6 0.0 53.8 54.6 60.9 94.4 0.0 78.8
Qwen2-VL [46] 7B 51.4 37.9 47.6 41.2 12.8 53.0 64.1 93.0 78.4 14.3 55.0 52.3 84.1 88.9 3.2 18.2
Qwen2-VL [46] 72B 59.9 42.4 51.2 50.8 17.4 52.0 71.2 97.7 84.5 6.4 77.5 66.1 76.8 98.2 16.1 84.9
Qwen2.5-VL [6] 7B 59.2 44.0 56.2 51.3 18.5 52.0 68.0 98.8 88.7 0.0 65.0 65.7 89.9 100.0 3.2 78.8
DeepSeek-VL2-Tiny [47] 3B 32.6 29.5 45.2 34.4 4.6 32.0 39.0 76.7 32.0 0.0 37.5 29.4 39.1 57.4 0.0 18.2
DeepSeek-VL2-Small [47] 16B 51.5 37.6 47.6 43.6 12.5 48.5 63.8 98.8 70.1 11.1 60.0 53.2 76.8 53.7 11.3 81.8
InternVL2 [11] 8B 48.4 31.9 44.3 38.0 0.0 48.5 62.9 98.8 62.9 4.8 70.0 50.5 68.1 75.9 0.0 66.7
InternVL2.5 [10] 8B 50.7 35.0 48.8 36.0 0.0 60.0 65.6 98.8 72.2 4.8 70.0 51.4 68.1 77.8 0.0 69.7
InternVL2.5 [10] 38B 63.1 44.0 59.9 52.0 2.5 66.0 78.8 98.8 92.8 38.1 72.5 66.5 98.6 96.3 3.2 69.7
Closed-Source Generic MLLMs
GPT-4o - 53.3 42.8 58.4 53.2 1.1 62.5 60.7 72.1 84.5 1.6 66.3 56.4 92.8 72.2 1.6 57.6
GPT-o1 - 36.5 15.8 33.2 11.6 0.0 14.0 41.4 75.6 52.6 0.0 23.8 52.3 82.6 81.5 0.0 39.4
Open-Source Mathematical MLLMs
Math-LLaVA [41] 13B 40.0 27.9 34.4 32.4 0.0 50.5 44.8 81.4 55.7 0.0 27.5 47.3 78.3 59.3 0.0 51.5
G-LLaVA [18] 7B 30.3 25.6 27.8 41.2 0.4 38.0 32.3 45.4 38.1 4.8 32.5 33.9 58.0 37.0 0.0 42.4
MultiMath [38] 7B 42.1 31.2 44.0 30.4 1.07 53.0 46.7 81.4 53.6 4.7 33.8 48.6 79.7 57.4 3.2 33.8
SVE-Math-DeepSeek [55] 7B 46.6 35.4 52.4 36.0 3.56 51.0 49.4 77.9 62.9 1.5 41.3 55.1 81.2 75.9 1.6 69.7
SVE-Math-DeepSeek+ (ours) 7B 68.4 84.6 75.8 88.4 82.9 97.5 54.1 85.3 65.8 20.3 45.0 60.7 85.1 78.4 1.6 75.7
2.3 Perception-oriented Training Dataset (GeoPeP)
Evaluation results on MATHGLANCE reveal that both open-source and closed-source MLLMs struggle to identify relevant visual regions in symbolic and abstract mathematical diagrams, despite their strong performance in theoretical reasoning and numerical computation. This contrasts with human cognitive abilities, where low-level perceptual tasks are typically solved rapidly compared to high-level reasoning tasks. To overcome such perceptual limitations, we construct a perception-oriented training dataset, GeoPeP, which enhances geometric understanding and visual grounding, ultimately benefiting complex reasoning. GeoPeP, comprising 200K high-quality image-text pairs in both diagram-caption and instruction-conversation formats, provides detailed geometric primitive information, including shapes, vertices, relationships, and fine-grained bounding box coordinates.
[…]
Overall, GeoPeP contributes to MLLM training by: (1) providing clear object attributes and their relationships, akin to graph nodes and edges, in training QA pairs; (2) offering fine-grained bounding box coordinates of elements, enabling models to systematically learn spatial awareness; and (3) integrating with reasoning-based CoT mathematical visual datasets during self-fine-tuning stage, allowing models to both perceive and reason accurately.
3 Experiments
In this section, we first describe the experimental setup (§3.1). We then present a comprehensive evaluation of 15 recent MLLMs (§3.2), demonstrating that while humans achieve high accuracy, MATHGLANCE remains challenging for existing models. Finally, we conduct comprehensive ablations and analyses, including detailed comparisons with math-specific MLLMs on both reasoning and perception benchmarks; investigation of key factors affecting MLLMs’ perception during inference; and evaluation of how GeoPeP improves both perception and reasoning (§3.3).
3.1 Experimental Setup
Generic and Mathematical MLLMs. We evaluate 15 MLLMs on MATHGLANCE across plane geometry, solid geometry, and graphs, including closed-source generic models such as GPT-4o and GPT-o1, as well as open-source models like LLaVA-v1.5 [32], mPLUG-Owl3 [48], InternLM-XComposer2 [16], Qwen2VL [46], Qwen2.5VL [6], DeepSeek-VL2 [47], InternVL2 [11], and InternVL2.5 [10]. Additionally, we assess math-specific MLLMs, including SVE-Math-DeepSeek [55], Math-LLaVA [41], G-LLaVA [18], and MultiMath [38]. This comprehensive evaluation provides insights into the diagram perception capabilities of state-of-the-art multimodal models.
Implementation Details. For open- and closed-source generic MLLMs, we follow the official inference settings, including temperature, number of beams, and maximum token length. For open-source mathematical LLMs, we adopt the standard configurations from SVE-Math-DeepSeek, setting the temperature to 0, the number of beams to 1, and the maximum token length to 1024. Notably, MultiMath [38] uses a temperature of 0.2; therefore, for a fair comparison, we rerun the inference with the temperature set to 0. We retain the original 512$\times$512 image resolution throughout all experiments. Choices are extracted using predefined rules tailored to each MLLM’s output format.
3.2 Main Results
The performance comparison of current MLLMs on plane geometry, solid geometry [24], and graphs [25] is summarized in Table 2. To assess models’ geometric diagram perception abilities, we evaluate their performance across shape classification, object counting, relationship identification, and object grounding. Key findings reveal that both generic and mathematical MLLMs face significant challenges in geometric perception tasks.
Generic MLLMs. General-purpose models trained on diverse datasets, including tables, charts, and documents [25, 9, 51], as well as visual grounding datasets [49, 40], still perform poorly on mathematical diagram perception. In particular, their performance on plane geometry remains low, with most models scoring below 45% on average. For solid geometry and graphs, general-purpose models significantly outperform mathematical MLLMs, due to their exposure to large-scale FigureQA [25], CLEVR [24], and various chart understanding datasets [51]. However, they still fail in fine-grained box-level tasks, with most models achieving 0 accuracy, and lag significantly behind human-level perception. These evaluation results highlight the limitations of generic MLLMs in mathematical diagram perception. Despite being trained on large-scale and diverse natural image datasets (e.g., over 2T tokens for Qwen2.5-VL and DeepSeek-VL2), these models struggle with the abstract nature of mathematical diagrams, which are defined by precise geometric structures and symbolic notations. Unlike natural images—where models can exploit semantic priors (e.g., people typically appear on the ground, birds in the sky)—mathematical diagrams lack such contextual cues. Their symbolic and structured nature requires genuine visual understanding rather than superficial pattern recognition. Without this, models fail to identify where to look, leading to poor perception performance. Notably, scaling up model size is neither an optimal nor an effective solution, as it provides only marginal gains in perception compared to reasoning benchmarks. For instance, increasing Qwen2VL from 7B to 72B improves top-1 accuracy by 22.3% on MathVista but only 8.3% on MATHGLANCE.
Mathematical MLLMs. Models designed for mathematical reasoning, such as MultiMath, Math-LLaVA, and SVE-Math-DeepSeek, exhibit strong reasoning performance on benchmarks like MathVerse [53], MathVista [34], and GeoQA [18]. They struggle with fundamental geometric perception, particularly fine-grained object localization, achieving near-zero accuracy at the default IoU threshold of 0.65. Since SVE-Math-DeepSeek incorporates a geometric visual encoder trained with detection and boundary segmentation losses, it outperforms other mathematical MLLMs on our benchmark across both cross-level and fine-level perception tasks. Furthermore, as it is not exposed to solid geometry diagrams or graphical images during MLLM training stages, it provides a more controlled experimental setting for analyzing perception capabilities compared to general-purpose models. Therefore, we select SVE-Math-DeepSeek as our base model to explore enhancements in both perception and reasoning. The comparison between SVE-Math-DeepSeek and other mathematical MLLMs is shown in Tab. 3. SVE-Math-DeepSeek trained on GeoPeP significantly improves solid geometry perception. Even without direct training on solid geometry and graphs, our model outperforms others due to its enhanced ability to discriminate relationships and understand spatial configurations. Low-level diagram structure understanding is essential for high-level reasoning skills, contributing to $\sim+4\%$ across three mathematical reasoning benchmarks. We leave the exploration of structure-aware training samples for solid geometry and graphs to future work.
Table 3: Performance comparison of mathematical MLLMs on both math perception and reasoning benchmarks.
Model MATHGLANCE MathVerse MathVista GeoQA
Plane Soild Graphs
Math-LLaVA [41] 27.9 44.8 47.3 20.1 46.6 60.7
G-LLaVA [18] 25.6 31.3 33.9 17.8 25.6 64.2
MultiMath [38] 31.2 45.7 48.6 25.9 49.3 74.1
SVE-Math-DeepSeek [55] 35.4 49.4 55.1 24.3 48.7 72.8
SVE-Math-DeepSeek+ 84.6 54.1 60.7 28.1 51.3 76.2
Figure 5: We evaluate five key factors affecting the perception ability of MLLMs: the number of obejcts (# obejct), visual quality, visual and textual distractors, and Chain-of-Thought (CoT) reasoning strategies (see Fig. 6 for details).
Figure 6: Analyzing perceptual errors: CoT vs. direct reasoning.
Figure 7: Response comparisons between SVE-Math-DeepSeek $\!\!{}^{+}$, SVE-Math-DeepSeek, and InterVL2.5.
3.3 Ablation Study
3.3.1 Key Factors Influencing Diagram Perception
We examine five key factors: the number of objects, object quality, visual distractors, text distractors, and Chain-of-Thought (CoT) responses. While those factors are by no means exhaustive, they aim to illuminate some of the fundamental perceptual limitations of current MLLMs, providing insights for practical applications and future improvements. The evaluation results on plane geometry from MATHGLANCE across nine MLLMs are presented in Fig.5, and response error patterns are shown in Fig.6.
The Number of Objects. The complexity of visual perception tasks increases with the number of objects in an image. We observe that as the number of geometric shapes grows, models struggle with accurate object localization, shape classification, and counting. Fig. 1 presents the overall results across all perception tasks for eight MLLMs and our method. When scaling the number of objects from 1 to $N$ ($N<5$), the Qwen-series model’s performance drops by approximately 12%, the DeepSeek-series by 18%, and GPT-4o by 15%. Our method and its baseline show declines of 6% and 13%.
Table 4: Accuracy (%) on plane geometry perception w.r.t. object quality (w/ or w/o Gaussian noise) and visual distractor injection.
MATHGLANCE Noise $\sim\mathcal{N}({0},{0.3})$ Distractors
w/o w/Δ↑↓ w/o w/Δ↑↓
Qwen2VL-7B 29.8 33.9↑4.1 29.8 26.2↓3.6
Qwen2.5VL-7B 33.9 32.1↓1.8 33.9 31.0↓2.9
DeepSeek-VL2-Tiny 27.9 23.8↓4.1 27.9 22.6↓5.3
DeepSeek-VL2-Small 27.9 37.5↑9.6 27.9 25.9↓2.0
LLaVa-1.5-13B 26.8 31.6↑4.8 26.8 29.1↑2.3
InternVL2.5-8B 25.6 20.8↓4.8 25.6 19.6↓6.0
GPT-4o 32.7 30.9↓1.8 32.7 28.6↓4.1
SVE-Math-DeepSeek-7B 32.8 28.0↓4.8 32.8 29.2↓3.6
SVE-Math-DeepSeek +-7B 80.7 78.3↓2.4 80.7 80.3↓0.4
Visual Distractors. We draw visual distractors—irregular scribbles, wedge-shaped angle markers, and auxiliary lines—to evaluate MLLMs’ ability to focus on relevant geometric elements (see Fig.9). Despite prompting models to ignore these distractors, most are still negatively affected. Close-up results are shown in Tab. 4. All evaluated MLLMs show a performance drop of 2–10% under visual distractors, except LLaVA-1.5-13B, which improves by 2.3%. Our model remains robust, showing minimal sensitivity to distractors and better visual focus on geometric primitives.
Object Quality. For visual fidelity analysis, we apply Gaussian noise to degrade object quality. As shown in Tab. 4, with a default variance of 0.3, most models show performance degradation, while Qwen2VL, DeepSeek-VL2-Small, and LLaVA-1.5-13B unexpectedly improve. This may stem from their exposure to visually degraded data during training, making noisy images more aligned with their learned distribution—especially in the absence of clean geometric diagrams in the training corpus. To enable a more comprehensive evaluation and mitigate data bias, we introduce large Gaussian noise with variances of 0.5/0.8 to assess model performance under severe visual distortion. As noise increases, all models struggle with geometric recognition, showing significant performance drops. Fig.10 reports perception accuracy under noise, and distorted examples are shown in Appendix Fig.11, where at variance 0.8, even humans fail to interpret the diagrams.
Table 5: Performance of relationship identification (rlat) on MATHGLANCE under different textual distractor settings.
MATHGLANCE (rlat) Unrela. infor. Conflicts
w/o w/Δ↑↓ w/o w/Δ↑↓
Qwen2VL-7B 53.0 48.5↓4.5 53.0 24.5↓28.5
Qwen2.5VL-7B 52.0 56.0↑4.0 52.0 30.0↓22.0
DeepSeek-VL2-Tiny 32.0 34.5↑2.5 32.0 3.5↓28.5
DeepSeek-VL2-Small 48.5 42.5↓6.0 48.5 20.5↓28.0
LLaVa-1.5-13B 42.0 37.5↓4.5 42.0 16.0↓26.0
InternVL2.5-8B 60.0 49.0↓11.0 60.0 18.0↓42.0
GPT-4o 62.5 63.5↑1.0 62.5 38.0↓24.5
SVE-Math-DeepSeek-7B 51.0 49.0↓2.0 51.0 15.0↓36.0
SVE-Math-DeepSeek +-7B 88.0 86.5↓1.5 88.0 74.0↓14.0
Text Distractors. Researchers have observed that large language models (LLMs) often dominate MLLMs, relying heavily on knowledge shortcuts rather than genuinely perceiving visual content. To investigate this, we evaluate two settings in the relationship identification task. The first involves introducing unrelated information w.r.t. the target task. For example, when the task is to identify a mathematical relationship, the model is additionally given irrelevant details about foreground/background colors, box shapes, or object attributes. The second setting introduces contradictory cues; for instance, if two lines are parallel, the model is explicitly prompted that they are perpendicular. The results in Tab. 5 show that providing conflicting knowledge significantly impairs the perceptual ability of all evaluated MLLMs. The impact is especially pronounced in models with weaker perception capabilities, such as InternVL2.5-8B (48.4% in avg.), while stronger perceptual models like Qwen2.5-VL-7B (59.2% in avg.) are less affected. Our method demonstrates greater robustness, though it still experiences a 14% performance drop. These findings highlight that reliance on knowledge shortcuts remains a critical issue, particularly in abstract diagram recognition, where models must genuinely understand visual content rather than rely on superficial pattern matching.
Chain-of-Thought (CoT) Response. CoT reasoning is designed to enhance step-by-step logical inference, yet its effectiveness in visual perception tasks remains uncertain. Our analysis shows that while CoT improves textual reasoning, it does not directly enhance spatial or geometric understanding. Models incorporating CoT reasoning often struggle with fundamental perception tasks, leading to significant CoT errors. Models often generate excessive yet irrelevant rationale misaligned with the diagram, ultimately resulting in incorrect responses, particularly in Qwen-series models and GPT-4o (over 30% in Fig. 6). See §G for CoT response examples in the Appendix.
Table 6: Performance comparison w.r.t. variants trained on GeoPeP and other mathematical visual alignment datasets. $\bigstar$ indicates the pretrained model continuously self-fine-tuned using MathV360K and instruction-formatted GeoPeP.
Model MATHGLANCE MathVerse MathVista
Plane Soild Graphs
SVE-Math-DeepSeek 35.4 49.4 55.1 24.3 48.7
Align. +AutoGeo ($\triangle$) 18.6 34.7 40.4 17.8 45.0
+MAVIS ($\Diamond$) 6.0 2.5 6.9 6.8 34.3
+GeoPeP ($\heartsuit$) 41.6 40.2 49.2 19.7 46.2
SFT $\triangle$ $\bullet$ $\bigstar$ 79.6 44.2 55.1 25.2 48.3
$\Diamond$ $\bullet$ $\bigstar$ 78.1 42.3 54.1 23.3 47.1
$\heartsuit$ $\bullet$ $\bigstar$ 84.6 54.1 60.7 28.1 51.3
3.3.2 Effect of GeoPeP
Building on SVE-Math-DeepSeek, we first train the projector while freezing the LLM and visual encoder during the visual-language alignment stage using either AutoGeo, MAVIS, or GeoPeP image-caption alignment datasets. As shown in Tab. 6, training with GeoPeP yields a 6.2% gain on perception tasks, but suffer a substantial drop on the other two datasets ($\sim$-27%). Using these three pretrained models, we then conduct self-fine-tuning (SFT) with MathV360K and instruction-formatted GeoPeP. The model trained with GeoPeP achieves a notable $\sim$4improvement on both MathVerse and MathVista compared to SVE-Math-DeepSeek, whereas other variants underperform on MathVista. We hope this observation encourages further research into more effective methods and training datasets to enhance the visual perception capabilities of mathematical MLLMs, as clearer visual understanding enables more faithful and accurate reasoning responses (Figs.7/15-16).
4 Conclusion
In this paper, we introduce MATHGLANCE, a perception-demanding benchmark designed to evaluate geometric perception in MLLMs. While these tasks appear trivial for humans to solve ‘at a glance’, they present significant challenges for current models. Unlike semantic-rich natural images, mathematical diagrams are structured and abstract, characterized by symbolic elements and their precise interrelationships. Thus, they require genuine visual interpretation rather than pattern recognition or textual shortcuts. Through extensive analysis, we identify key factors affecting diagam perception, highlighting the importance of high-quality and structure-aware training datasets to enable MLLMs to both see and reason effectively. Our findings suggest that improving low-level geometric perception is essential for advancing high-level reasoning, bridging the gap between visual understanding and logical inference.
Appendix
Appendix A Related Work
A.1 Mathematical Reasoning Benchmark
To evaluate MLLMs performance across different domains, various benchmarks [26, 50, 27, 19, 33] have been proposed, primarily focusing on natural scene understanding. However, benchmarks specifically designed for multimodal mathematical reasoning remain scarce.
Early benchmarks such as MathQA [4], UniGeo [8]Geometry3k [35], GEOS [39], and GeoQA++ [5] introduced multimodal mathematical tasks but were limited in scope, often focusing on specific subdomains like plane geometry. More recent efforts have sought to provide broader and more diverse evaluations. MMMU [52] assesses multimodal mathematical understanding with an emphasis on symbolic reasoning and word problem-solving. MathVista [34] targets geometry-related tasks by integrating both real-world and synthetic diagrams to evaluate visual reasoning. MathVerse [53] expands this by incorporating a wider range of multimodal challenges involving charts, graphs, and structured visual content. MATH-V [45] further addresses the limitations of existing benchmarks by curating 3,040 high-quality math problems sourced from real-world competitions. While these benchmarks introduce visual elements, their core focus remains on assessing mathematical reasoning. It remains unclear whether the performance on these tasks truly reflects the models’ ability to comprehensively understand the symbolic information in diagrams. The ability to accurately interpret mathematical symbols, diagrams, and spatial structures is a fundamental component of solving multimodal math problems, yet current benchmarks place limited emphasis on this aspect.
A.2 MLLMs for Math
While multimodal large language models [3, 32, 42] have made significant strides in vision-language understanding, their mathematical reasoning capabilities remain limited. Foundation models such as GPT-4V [2], Qwen2-VL [46], and Deepseek-VL2 [47] perform well on general multimodal tasks but struggle with mathematical symbol recognition, spatial reasoning, and logical deduction, making them inadequate for vision-based math problem-solving.
To address these limitations, recent efforts have introduced math-specific MLLMs for improving mathematical reasoning. AlphaGeometry [43] achieves state-of-the-art results in geometry by leveraging theorem-proving and reinforcement learning. However, it relies solely on text-based diagram descriptions, lacking direct image processing. G-LLaVA [18] extends LLaVA with geometric reasoning capabilities but struggles with complex visual structures and generalization beyond plane geometry. UniMath [29] integrates structured math representations for solving visual word problems but remains focused on symbolic reasoning, limiting its handling of free-form mathematical diagrams. MatCha [31] specializes in chart-based reasoning, extracting quantitative relationships from structured visual data. However, its reliance on predefined formats limits adaptability to unstructured mathematical visuals.
Beyond architecture improvements, MAVIS [54] introduces an automated data engine to generate large-scale mathematical visual datasets, reducing annotation costs while ensuring high-quality diagram-caption pairs and problem-solving rationales. However, its diagrams are constructed by simply combining basic geometric shapes, lacking mathematical relationship constraints such as perpendicularity and parallelism. AutoGeo [21] considers special properties of lines, such as midlines and radii, as foundational to many geometric theorems, incorporating these properties into geometric figures. Reverse Chain-of-Thought (R-CoT) [15] introduces the Geometry Generation Chain to generate the geometry image and corresponding description. However, current data engines still fail to explore the underlying structures in mathematical diagrams, leading to ambiguous captions and redundant information that negatively impact MLLMs’ perception abilities. Consequently, models trained on such datasets perform poorly on perception-demanding tasks, even falling below random guessing on multiple-choice questions. Thus, building a mathematical visual dataset that enables models to perceive and reason is a urgent task.
Appendix B Visualization of MATHGLANCE
MATHGLANCE is specifically designed to evaluate perception-demanding tasks in Multimodal Large Language Models (MLLMs), focusing on four core visual understanding capabilities: shape classification, object counting, object grounding, and relationship identification. These tasks, commonly studied in classical computer vision, are typically solvable by humans with minimal cognitive effort. Our benchmark spans three key domains—plane geometry, solid geometry, and mathematical graphs—to ensure broad coverage of visual representations encountered in educational and scientific contexts. Fig. 8 provides illustrative examples from MATHGLANCE, demonstrating the diagrammatic input and corresponding question–answer (Q&A) pairs used for evaluation.
Figure 8: Visualization of sample cases from MATHGLANCE. (a), (b), and (c) correspond to problems related to Plane Geometry, Solid Geometry, and Graphs, respectively.
Appendix C Problem Templates
This section introduces the problem templates employed in MATHGLANCE, with illustrative examples provided in Tab. 7. We adopt a template-based generation engine to systematically construct diverse perception-oriented Q&A pairs. Each question is generated by parsing the structured JSON annotations of a given image, which include information on shapes, object attributes, spatial positions, and inter-object relationships. These elements are then filled into carefully designed templates (Tab. 7) to produce grammatically correct and semantically meaningful Q&A pairs.
Our template designs are tailored to accommodate diverse subject domains—including plane geometry, solid geometry, and graph-based diagrams—and span a spectrum of perceptual complexity, from coarse-level to fine-grained tasks. For example, a coarse-level shape classification template may pose a question such as, “What is the shape of the object with {vertices}?”, where the correct answer is directly retrieved from the structured annotations. For fine-grained object localization, we adopt a grounding template similar to that used in [47], prompting models with: “Please provide the bounding box coordinates of the region this sentence describes: {shape} {vertices}”. This allows us to evaluate the model’s ability to extract precise spatial information and align language with visual primitives at a granular level.
We design three types of questions: multiple-choice, true/false, and open-ended. Specifclly, for multiple-choice questions, we incorporate plausible distractors to challenge the model’s geometric perception. These distractors are selected from geometric candidate pools, varying from visually similar to clearly distinguishable options, ensuring a balanced evaluation of fine-grained visual discrimination. By leveraging this template engine, MATHGLANCE ensures consistent question formulation, controlled variation in difficulty, and robust coverage of geometric primitives and their relationships, forming a reliable testbed for evaluating diagram perception in MLLMs.
Figure 9: Visualizing distractors w.r.t. adding Gaussian noise, drawing irregular scribbles and incorporating wedge-shaped symbols or auxiliary lines.
Figure 10: Perceptual perfromance w.r.t. varying Gaussian noise variance.
Figure 11: Visualization of distorted images under varying levels of Gaussian noise.
Appendix D Reformat Dataset for Solid Geometry & Graphs
Solid Geometry. CLEVR [24] is a synthetic Visual Question Answering (VQA) dataset containing 3D-rendered objects. Each object in the scene is defined by its 3D position and a set of attributes, including size (small or large), shape (cube, cylinder, sphere), material (rubber or metal), and color (gray, blue, brown, yellow, red, green, purple, cyan). However, CLEVR’s annotation format is incompatible with our template-based pipeline for generating perception evaluation QA pairs. To resolve this, we reformat the original annotations into structured JSON, following the same format as the synthetic data process. Specifically, we extract object information from CLEVR’s scene annotations and organize them into shape information and object attributes. We then leverage an object’s 3D location and its direction for calculating its 2D top-left and bottom-right coordinates, from which relative spatial relationships are derived.
Once the reformatted JSON annotations are generated, we apply the same template-based QA generation process as designed for plane geometry, with an additional uniqueness check for shape classification. To ensure clear and consistent labeling, we filter out ambiguous cases where identical attribute combinations correspond to different shapes within the same image. For example, if an image contains multiple distinct shapes with the same attributes (e.g., a large rubber blue square and a large rubber blue cylinder), we exclude classification questions that rely on the attribute combination ‘large rubber blue’ to prevent ambiguity.
Graphs. FigureQA [25] is a visual reasoning dataset containing over one million question-answer pairs, grounded in synthetic, scientific-style figures, including line plots, dot-line plots, vertical and horizontal bar graphs, and pie charts. The official annotations capture various relationships between plot elements and evaluate characteristics such as maximum, minimum, smoothness, and intersection, all framed as binary yes/no questions. Additionally, FigureQA provides numerical data annotations used to generate each figure, along with bounding-box annotations for all plot elements. Similar to CLEVR, we reformat FigureQA annotations compatible with template-based pipelines.
Specifically, the shapes include five graph types as defined in the official specification, with each element’s attribute represented by its unique color. Additionally, we store the bounding box coordinates of each foreground element (such as lines, bars, and pie slices), along with legends and titles, as ground truth for the grounding task. For relationships, we follow the binary true/false question-answering problems to evaluate the model’s understanding of geometric and graphical relationships.
Appendix E Example Illustration of GeoPeP
Figs. 12-14 demonstrate how GeoPeP delivers explicit geometric information—covering object count, shape classification, fine-grained bounding box coordinates, and inter-object relationships—presented in both caption-style and instruction-following conversational formats.
Appendix F Visual Distractors
To evaluate the robustness of MLLMs’ perceptual capabilities, we introduce a set of visual perturbations through data augmentation. These include Gaussian noise, irregular scribbles, wedge-shaped symbols, and auxiliary lines, as shown in Fig. 9. The aim is to simulate real-world visual ambiguities and assess whether MLLMs can retain accurate geometric understanding under degraded conditions. These controlled distortions provide a rigorous benchmark for evaluating the models’ ability to extract and interpret mathematical structures from visually complex inputs.
By gradually increasing the Gaussian noise level, we can systematically evaluate its impact on the ability of MLLM to recognize mathematical structures. As shown in Fig. 11, the image transitions from a noise level of 0.1 to 0.8, progressively blurring the geometric features. We observe that as the noise intensity increases, the difficulty of recognizing mathematical structures also rises, posing greater challenges to accurate recognition and reasoning. When the noise level reaches $\sim\mathcal{N}(\mathbf{0},\mathbf{0.8})$, the mathematical structures become nearly indistinguishable to the human eye, resulting in a significant performance drop on MATHGLANCE, as illustrated in Fig. 10.
Appendix G Case Study
Model Responses. In Fig. 15 and Fig. 16, we present a comparative analysis of the performance of different Multimodal Large Language Models (MLLMs), including SVE-Math-DeepSeek [55] (baseline), SVE-Math-DeepSeek+, and InterVL2.5 [10]. Our evaluation reveals that SVE-Math-DeepSeek+ achieves a significant improvement in response accuracy compared to its baseline, SVE-Math-DeepSeek. This improvement suggests that our carefully curated training data effectively enhances the model’s mathematical perception, enabling it to better comprehend problem structures, reason through mathematical concepts, and generate more precise answers. Furthermore, we observe that even InterVL2.5, a strong competitor in the field, produces incorrect responses in certain cases. Upon closer examination, these errors often stem from inaccurate mathematical perception rather than purely computational mistakes. This observation reinforces the notion that an MLLM’s ability to accurately perceive and interpret mathematical content is a critical factor in achieving high performance. Overall, our findings highlight the fundamental role of precise perception in mathematical reasoning, akin to its importance in vision and language understanding. Just as accurate perception is essential for tasks such as image recognition and natural language processing, it is equally vital for MATH-related problem-solving. Our study underscores that enhancing a model’s mathematical perception can lead to substantial gains in accuracy and reliability.
Error Examples. In this section, we provide more detailed error examples of GPT-4o, Qwen2.5-VL-7B, and our 7B model. We categorize the errors into two types: Chain-of-Thought (CoT) Errors and Recognition Errors. CoT errors occur when the model engages in step-by-step reasoning for perception questions but ultimately provides incorrect answers. Recognition errors, on the other hand, arise when the model attempts direct answering without reasoning yet fails to produce the correct result. Representative examples for each error type are illustrated in Figs. 17-26.
Figure 12: Sampels of caption-style and instruction-following GeoPeP.
Figure 13: Sampels of caption-style and instruction-following GeoPeP.
Figure 14: Sampels of caption-style and instruction-following GeoPeP.
Table 7: Examples of problem templates used by MATHGLANCE on different source data acorss different tasks.
Source Task Three randomly chosen examples from hundreds.
Plane Geometry cls
What is the shape of object {vertices} in the image? Choices: A:{a} B:{b} C:{c} D:{d}.
Can you identify the type of object {vertices} in the picture? Choices: A:{a} B:{b} C:{c} D:{d}.
Can you identify the type of object in the picture? Choices: A:{a} B:{b} C:{c} D:{d}.
What is the shape of the object in the image? Choices: A:{a} B:{b} C:{c} D:{d}.
cnt
There {be} {num} {shape} in the picture, right? Choices: A:Yes B:No.
{be} there {num} {shape} in the picture? Choices: A:Yes B:No.
You can see {num} objects in the picture, can’t you? Choices: A:Yes B:No.
There should be {num} shapes in the picture, correct? Choices: A:Yes B:No.
How many {shape}s can you find in the picture? Choices: A:{a} B:{b} C:{c} D:{d}.
Please count all the {shape}s in the image. Choices: A:{a} B:{b} C:{c} D:{d}.
What is the total number of shapes in the picture? Choices: A:{a} B:{b} C:{c} D:{d}.
How many objects can you identify in the image? Choices: A:{a} B:{b} C:{c} D:{d}.
Please identify and select all the {shape}s in the picture. Choices: A:{a} B:{b} C:{c} D:{d}.
Find and select all the {shape}s in the picture. Choices: A:{a} B:{b} C:{c} D:{d}.
grd Please provide the bounding box coordinate of the region this sentence describes: {shape} {vertices}.
rlat
Can the relationship {preposition} {shape} in the image be described as ”{relation}”? Choices: A:Yes B:No.
Does the image show the relationship {preposition} {shape} as ”{relation}”? Choices: A:Yes B:No.
Is {shape1} described as being in the ’{relation}’ relative to {shape2} in the image? Choices: A:Yes B:No.
Is {shape1} said to be in the ’{relation}’ relative to {shape2} in the image? Choices: A:Yes B:No.
What is the relationship {preposition} {shape} in the picture? Choices: A:{a} B:{b} C:{c} D:{d}.
Can you identify the relationship {preposition} {shape} in the image? Choices: A:{a} B:{b} C:{c} D:{d}.
What is the relative position of {shape1} to {shape2} in the image? Choices: A:{a} B:{b} C:{c} D:{d}.
What is the spatial relationship of shape1 to shape2 in the image? Choices: A:{a} B:{b} C:{c} D:{d}.
Soild Geometry cls
What is the shape of the {size} {color} object made of {material} in the image? Choices: A:{a} B:{b} C:{c} D:{d}.
Can you identify the type of {size} {color} object with {material} material in the picture? Choices: A:{a} B:{b} C:{c} D:{d}.
cnt
There {be} {num} {shape} in the picture, right? Choices: A:Yes B:No.
{be} there {num} {shape} in the picture? Choices: A:Yes B:No.
There are {num} objects in the picture, right? Choices: A:Yes B:No.
Are there {num} shapes in the picture? Choices: A:Yes B:No.
How many {shape}s can you find in the picture? Choices: A:{a} B:{b} C:{c} D:{d}.
Please count all the {shape}s in the image. Choices: A:{a} B:{b} C:{c} D:{d}.
Count the shapes in the image. Choices: A:{a} B:{b} C:{c} D:{d}.
How many shapes can you visually identify in the image? Choices: A:{a} B:{b} C:{c} D:{d}.
How many {size} {color} {material} objects are there in the image? Choices: A:{a} B:{b} C:{c} D:{d}.
How many {size} {color} {material} objects are present in the image? Choices: A:{a} B:{b} C:{c} D:{d}.
grd Please provide the bounding box coordinate of the region this sentence describes: {size} {color} {material} {shape}.
rlat
Is it correct that {shape1} is described as being in the ’{relation}’ relative to {shape2} in the image? Choices: A:Yes B:No.
Is {shape1} described as being in the ’{relation}’ relative to {shape2} in the image? Choices: A:Yes B:No.
Is {shape1} said to be in the ’{relation}’ relative to {shape2} in the image? Choices: A:Yes B:No.
Can you confirm that {shape1} is in the ’{relation}’ relative to {shape2} in the image? Choices: A:Yes B:No.
In the image, where is {shape1} in relation to {shape2}? Choices: A:{a} B:{b} C:{c} D:{d}.
What is the relative position of {shape1} to {shape2} in the image? Choices: A:{a} B:{b} C:{c} D:{d}.
What is the spatial relationship of {shape1} to {shape2} in the image? Choices: A:{a} B:{b} C:{c} D:{d}.
Describe how {shape1} is situated relative to {shape2} in the image. Choices: A:{a} B:{b} C:{c} D:{d}.
Graphs cls
What type of chart is shown in the image? Choices: A:{a} B:{b} C:{c} D:{d}.
Identify the type of chart in the image. Choices: A:{a} B:{b} C:{c} D:{d}.
Which of the following best describes the chart graph in the image? Choices: A:{a} B:{b} C:{c} D:{d}.
Which kind of chart does the image represent? Choices: A:{a} B:{b} C:{c} D:{d}.
What kind of data visualization is shown in the image? Choices: A:{a} B:{b} C:{c} D:{d}.
Can you identify the type of chart in the image? Choices: A:{a} B:{b} C:{c} D:{d}.
cnt
{be} there {num} {shape} in the picture? Choices: A:Yes B:No.
The picture contains {num} {shape}, each represented by a different color, doesn’t it? Choices: A:Yes B:No.
You can see {num} {shape} with different colors in the picture, can’t you? Choices: A:Yes B:No.
How many {shape}s are there in the image? Choices: A:{a} B:{b} C:{c} D:{d}.
Can you count the total number of {shape}s in the image? Choices: A:{a} B:{b} C:{c} D:{d}.
What is the total number of {shape}s in the picture? Choices: A:{a} B:{b} C:{c} D:{d}.
grd Please provide the bounding box coordinate of the region this sentence describes: {color} {shape}.
rlat {question} Choices: A:Yes B:No.
Figure 15: Response comparisons between SVE-Math-DeepSeek $\!\!{}^{+}$, SVE-Math-DeepSeek, and InterVL2.5.
Figure 16: Response comparisons between SVE-Math-DeepSeek $\!\!{}^{+}$, SVE-Math-DeepSeek, and InterVL2.5.
Figure 17: Examples of Chain-of-Thought (CoT) Errors in GPT-4o.
Figure 18: An example of Chain-of-Thought (CoT) Errors in GPT-4o.
Figure 19: An example of Chain-of-Thought (CoT) Errors in Qwen2.5-VL-7B.
Figure 20: An example of Chain-of-Thought (CoT) Errors in Qwen2.5-VL-7B.
Figure 21: An example of Chain-of-Thought (CoT) Errors in Qwen2.5-VL-7B.
Figure 22: Examples of Recognition Errors in GPT-4o.
Figure 23: Examples of Recognition Errors in GPT-4o (top) and Qwen2.5-VL-7B (bottom).
Figure 24: Examples of Recognition Errors in Qwen2.5-VL-7B.
Figure 25: Examples of Recognition Errors in SVE-Math-DeepSeek +-7B.
Figure 26: An example of Recognition Errors in SVE-Math-DeepSeek +-7B.